OpenStudy (anonymous):

how many different combinations of 6 volleyball players can be chosen from a pool of 15 players?

OpenStudy (anonymous):

15C6

jimthompson5910 (jim_thompson5910):

Order doesn't matter, so use the combination formula nCr = (n!)/(r!(n-r)!) 15C6 = (15!)/(6!*(15-6)!) 15C6 = (15!)/(6!*9!) 15C6 = (15*14*13*12*11*10*9!)/(6!*9!) 15C6 = (15*14*13*12*11*10)/(6!) 15C6 = (15*14*13*12*11*10)/(6*5*4*3*2*1) 15C6 = (3603600)/(720) 15C6 = 5005 So there are 5005 different combinations.